# Action-slot: Visual Action-centric Representations for Multi-label Atomic Activity Recognition in Traffic Scenes

### v1

@@ Abstract
In this paper, we study multi-label atomic activity recognition. Despite the notable progress in action recognition, it is still challenging to recognize atomic activities due to a deficiency in holistic understanding of both multiple road users’ motions and their contextual information. In this paper, we introduce Action-slot, a slot attention-based approach that learns visual action-centric representations, capturing both motion and contextual information. Our key idea is to design action slots that are capable of paying attention to regions where atomic activities occur, without the need for explicit perception guidance. To further enhance slot attention, we introduce a background slot that competes with action slots, aiding the training process in avoiding unnecessary focus on background regions devoid of activities. Yet, the imbalanced class distribution in the existing dataset hampers the assessment of rare activities. To address the limitation, we collect a synthetic dataset called TACO, which is four times larger than OATS and features a balanced distribution of atomic activities. To validate the effectiveness of our method, we conduct comprehensive experiments and ablation studies against various action recognition baselines. We also show that the performance of multi-label atomic activity recognition on real-world datasets can be improved by pretraining representations on TACO. We will release our source code and dataset. See the videos of visualization on the project page: https://hcis-lab.github.io/Action-slot/
@@
Figure 2: The distribution of atomic activity classes in our TACO dataset compared to the OATS dataset. Please note that, for space considerations, we omit the road topology notations of C and Z on the x-axis of the figure.
@@ 1 Introduction
In this paper, we focus on multi-label atomic activity recognition, a new task proposed by the OATS dataset for interactive traffic scenario understanding [3]. An atomic activity is a higher-level semantic motion pattern rooted in the underlying road topology. For instance, a group of vehicles turning left is described as Z1-Z4: C+, as shown in Figure Action-slot: Visual Action-centric Representations for Multi-label Atomic Activity Recognition in Traffic Scenes. The notions Z1 and Z4 are roadways within an intersection. A left-turn motion pattern is represented as Z1-Z4. We use C+ to denote a group of vehicles. This task presents distinctive challenges as it demands a comprehensive understanding of contextual information and motions of road users. Moreover, it requires the ability to break down activities within crowded and intricate traffic scenes from videos.
We identify the limitations of current action recognition models, showing less desirable performance in the OATS benchmark [3]. We argue that video-level representations, employing 3D ConvNet [12, 72, 29, 28] or transformers [27, 71] models, pose a challenge for classifiers in distinguishing individual atomic activities within traffic scenes due to the complicated structure of traffic scenes. On the other hand, object-aware representations [6, 75, 3] using object features [35] struggle to associate the relationship between the road structures and motions of objects. Then the question we like to answer in this paper: Can we learn visual representations that decompose multiple atomic activities from videos, without using object proposals?
@@
To address this limitation, we introduce the Traffic Activity Recognition (TACO) dataset, an extensive dataset for atomic activity recognition. We utilize the CARLA simulator [23] to gather instances of all conceivable activity classes, ensuring a well-balanced distribution, as illustrated in Figure 2. We again benchmark all methods on the TACO dataset and show that Action-slot outperforms all baselines by a large margin.
@@
    We provide qualitative evidence demonstrating that the learned action-centric representations can reliably identify distinct atomic activities in the visual domain, without explicit perception guidance.
@@ 3.1 Preliminary: Slot Attention
The Slot Attention module [48] can be viewed as a clustering algorithm that maps image patches input to a set of $K$ output slots $S\in\mathbb{R}^{K\times D_{\mathtt{slot}}}$, where $D_{\mathtt{slot}}$ is the dimension of each slot. Specifically, the slots $S$ are first initialized by randomly sampling $K$ vectors from a Gaussian distribution, where the parameter $K$ is usually defined as the maximum number of objects in an image. An input frame, i.e., image features $F\in$ $\mathbb{R}^{H\times W\times D_{\mathtt{in}}}$ with size $H\times W$ and dimension $D_{\mathtt{in}}$, is first flattened to patch tokens $F^{{}^{\prime}}\in$ $N\times D_{\mathtt{in}}$, where $N=H\times W$. Then the tokens are mapped to the slots using the dot product attention module. That is, the attention weight can be calculated as $A=\frac{1}{\sqrt{D}}k(F^{{}^{\prime}})\cdot q(S)\in\mathbb{R}^{N\times K}$, and $q$ and $k$ are linear transformations that map the input $F$ and slots $S$ to a common dimension $D$.
@@ 3.3 Action-slot
In this section, we describe the details of our design for learning action-centric representations. We follow [5] to use learnable parameters for slot initializations. We extract image features $F\in\mathbb{R}^{T\times H\times W}$ from the convolution-based encoder [34, 12, 29, 28] for each frame in a video clip and then flatten them to $F^{{}^{\prime}}\in\mathbb{R}^{N\times D_{\mathtt{in}}}$, where $N=T\times H\times W$. We add learnable 3D spatial-temporal positional embeddings $E\in\mathbb{R}^{T\times H\times W}$ to the tokens $F^{{}^{\prime}}$ [7, 4, 27, 5, 71].
@@
Instead of providing attention guidance for positive classes, which is non-trivial as explained previously, we introduce a regularization term to discourage action slots associated with negative classes (i.e., those activities not present in a video), from attending to any regions. We construct an attention mask $M_{\mathtt{neg}}$ wherein all element values are set to 0. We design a loss function: $L_{\mathtt{neg}}=\sum_{\{c|y_{c}=0\}}\textnormal{BCE}(\bar{A}_{c},M_{\mathtt{neg}})$, where $\bar{A}_{c}$ is the normalized attention output of slot $S_{c}$ for negative class $c$. By regularizing the action slots with negative classes, the competition mechanism makes slots with positive classes more competitive in the important regions. Note that the regularization does not require additional annotations. Finally, the full objective function of Action-slot is: $L_{\mathtt{all}}=L_{\mathtt{act}}+w_{\mathtt{bg}}L_{\mathtt{bg}}+w_{\mathtt{neg}}L_{\mathtt{neg}},$ where $w_{\mathtt{bg}}$ and $w_{\mathtt{neg}}$ are the weights to balance two terms (set as 0.5 and 1, respectively in this paper).
@@ 4.1 Datasets
TACO. We construct the TACO dataset in the CARLA simulator [23] to attain a balanced distribution encompassing 64 classes of traffic atomic activities. The TACO dataset comprises 5178 clips. Among them, 1148 are used for testing. We collect image and instance segmentation data with a size of 512 $\times$ 1536 pixels, and subsequently downsample each frame to 256 $\times$ 768 as input to the models.
@@ 4.2 Baselines
Object-aware models. Object-aware models [6, 75, 3] first extract object features via the RoI-Align operation [35]. To learn the relations between objects, ORN [6], ARG [75], and OATS [3] uses MLP, Graph Convolutional Network (GCN) [43], and spatical-temporal GCN [53] to enhance object features. OATS also uses an additional graph branch to encode low-level object trajectories. To extract trajectories of objects, we use Deep OC-SORT [51], the state-of-the-art model on the MOT20 benchmark [20] in terms of HOTA [50]. We follow OATS by setting the number of tracklets as 20 for the input of the models. To encourage the object-aware models [6, 75] to learn the context information, we concatenate object features with the global features, which are obtained by convolving the $5^{\mathtt{th}}$ ResBlock of X3D [28] with a 3D kernel size of 1. These models then aggregate all frames’ object features by summing them channel-wise, followed by a linear layer to output multi-class predictions for each proposal, with a channel number of $N_{cl}$ plus one negative class. Note that we do not implement the model proposed in OATS [3] due to the absence of source code.
Slot-based models. Slot-based models use the last ConvBlock of the encoders [34, 28] as backbone features. Note that we remove the last projection layer of X3D, which makes the size of some slot-based models slightly smaller than X3D. We re-implement existing slot-attention-based methods for videos [42, 25, 5, 82] to use the proposed “allocated slots” for multi-label atomic activity recognition. Specifically, SAVi [42] randomly initialize slots for each forward pass from a learned distribution, while MO [5] and Slot-VPS [82] treat slots as learnable queries [11]. For the slot update mechanism, SAVi and MO recurrently update slots in time, while Slot-VPS uses an extra self-attention [73] to update slots across frames. We then use their slots in the last frame as the input to the classifier.
@@
Method Backbone Pretrain S1 S2 S3 mAP
CSN [72, 3] ResNet152 IG65M 12.1 12.6 12.9 12.5
TPN [79, 3] ResNet50 ImageNet 11.6 13.3 12.9 12.6
SlowOnly [29, 3] ResNet50 ImageNet 11.2 14.7 12.9 13.0
SlowFast [29, 3] ResNet50 None 10.8 15.1 14.5 13.5
I3D (NL) [74, 3] ResNet50 ImageNet 11.9 15.5 14.0 13.8
I3D [12, 3] ResNet50 ImageNet 11.8 14.3 16.8 14.3
ORN [6, 3] ResNet50 ImageNet 16.8 13.4 18.1 16.1
ARG [75, 3] Inceptionv3 ImageNet 20.2 21.3 19.3 20.3
OATS [3] Inceptionv3 ImageNet 24.3 28.6 27.2 26.7
@@
Action-Slot (Ours) ResNet50 ImageNet 26.6 28.6 30.8 28.6
I3D [12] ResNet50 Kinetics-400 21.7 24.6 24.4 23.6
X3D [28] N/A Kinetics-400 30.4 33.2 30.6 31.4
CSN [72] ResNet101 Kinetics-400 43.1 47.1 44.3 44.8
SlowFast [29] ResNet50 Kinetics-400 36.1 36.6 34.2 35.6
ORN [6] X3D Kinetics-400 19.3 22.5 23.6 21.8
ARG [75] X3D Kinetics-400 24.8 25.9 29.3 26.7
@@
Action-Slot (Ours) X3D Kinetics-400 48.1 47.7 48.8 48.2
@@ 4.3 Implementation Details
Table 5: Results of various backbones for Action-slot on TACO.
@@
Table 5: Results of various backbones for Action-slot on TACO.
@@ 4.4 Main Results
OATS. In Table 1, we show experimental comparisons with video-level-based, object-aware, and slot-attention-based methods. In the top group, we use the ImageNet pre-trained ResNet50, following the evaluation protocol designed in OATS [3] We further compare all methods with advanced visual backbones pretrained from Kinetics-400 [41] and report the corresponding results in the bottom group. In both ImageNet and Kinetics pretrained settings, Action-slot achieves state-of-the-art performance on the OATS dataset. Specifically, Action-slot improves the performance of atomic activity recognition by identifying the region of interest via the proposed slot attention mechanism, without explicit perception guidance.
@@ 4.5 Ablation Study
Figure 6: Visualization of attention maps learned from TACO. Colored masks are from the allocated slots of activity Z4-Z1:B and Z1-Z2:K+. Note that Z4-Z1:K appears twice in the 1st and 5th frames independently.
@@ Appendix A Construction of The TACO Dataset
We collect scenarios with the following two approaches:
@@
    Scenario Runner [2]. Scenario runner, a scenario collection tool developed by CARLA [23], can explicitly generate a scripted scenario by defining a set of routes for road users. Although auto-pilot can help collect diverse traffic activities, it is difficult to generate some atomic activities frequently. For example, Z3-Z2:K+, a group of bicyclists turns left from the opposite roadway of ego-vehicle. Therefore, we leverage scenario runner to explicitly collect the scenarios that are difficult for the auto-pilot method. A scenario starts to collect when ego vehicle reaches a trigger point (i.e., location) predefined in the script. A scenario ends when all scripted actions are accomplished. To further enhance the diversity of scenarios, we randomly spawn road users surrounding the ego vehicle. Note that road users are not guaranteed to be involved in an atomic activity.
For both collection methods, we randomly set the weather and light conditions. Note that we exclude night scenes because of the poor visibility.
@@ Appendix C More Ablation Study of Action-slot
Object-guided v.s. Action-slot.
@@
Cross-Attention v.s. Slot Attention.

### v2

@@ Abstract
In this paper, we study multi-label atomic activity recognition. Despite the notable progress in action recognition, it is still challenging to recognize atomic activities due to a deficiency in holistic understanding of both multiple road users’ motions and their contextual information. In this paper, we introduce Action-slot, a slot attention-based approach that learns visual action-centric representations, capturing both motion and contextual information. Our key idea is to design action slots that are capable of paying attention to regions where atomic activities occur, without the need for explicit perception guidance. To further enhance slot attention, we introduce a background slot that competes with action slots, aiding the training process in avoiding unnecessary focus on background regions devoid of activities. Yet, the imbalanced class distribution in the existing dataset hampers the assessment of rare activities. To address the limitation, we collect a synthetic dataset called TACO, which is four times larger than OATS and features a balanced distribution of atomic activities. To validate the effectiveness of our method, we conduct comprehensive experiments and ablation studies against various action recognition baselines. We also show that the performance of multi-label atomic activity recognition on real-world datasets can be improved by pretraining representations on TACO. Our source code, dataset, and visualization videos are available at https://hcis-lab.github.io/Action-slot/
@@
Figure 2: The distribution of atomic activity classes in our TACO dataset compared to the OATS dataset. Please note that, for space considerations, we omit the road topology notations of corner (C) and roadway (Z) on the x-axis of the figure.
@@ 1 Introduction
In this paper, we focus on multi-label atomic activity recognition, a new task proposed by the OATS dataset for interactive traffic scenario understanding [3]. An atomic activity is a higher-level semantic motion pattern rooted in the underlying road topology. For instance, a group of vehicles turning left is described as Z1-Z4: C+, as shown in Figure 1. The notions Z1 and Z4 are roadways within an intersection. The left-turn motion pattern is represented as Z1-Z4. We use C+ to denote a group of vehicles. The combination of such motion patterns and road user types can compose 64 classes of atomic activities. This task presents distinctive challenges as it demands a comprehensive understanding of contextual information and motions of road users. Moreover, it requires the ability to break down activities within crowded and intricate traffic scenes from videos.
We identify the limitations of current action recognition models, showing less desirable performance in the OATS benchmark [3]. We argue that video-level representations, simply employing 3D ConvNet [12, 71, 28, 27] or transformers [26, 70] models, pose a challenge for classifiers in distinguishing individual atomic activities within traffic scenes due to the complicated structure of traffic scenes. On the other hand, object-aware representations [6, 74, 3] using object features [34] struggle to associate the relationship between the road structures and motions of objects. Then the question we like to answer in this paper: Can we learn visual representations that decompose multiple atomic activities from videos, without using object proposals?
@@
To address this limitation, we introduce the Traffic Activity Recognition (TACO) dataset, an extensive dataset for atomic activity recognition. We utilize the CARLA simulator [22] to gather instances of all conceivable activity classes, ensuring a well-balanced distribution, as illustrated in Figure 2. We again benchmark all methods on TACO and show that Action-slot outperforms all baselines by a large margin.
@@
    We provide qualitative evidence demonstrating that the learned action-centric representations can reliably identify distinct atomic activities in the visual domain without perception modules.
@@ 3.1 Preliminary: Slot Attention
The Slot Attention module [48] can be viewed as a clustering algorithm that maps image patches input to a set of $K$ output slots $S\in\mathbb{R}^{K\times D_{\mathtt{slot}}}$, where $D_{\mathtt{slot}}$ is the dimension of each slot. Specifically, the slots $S$ are first initialized by randomly sampling $K$ vectors from a Gaussian distribution, where the parameter $K$ is usually defined as the maximum number of objects in an image. An input frame, i.e., image features $F\in$ $\mathbb{R}^{H\times W\times D_{\mathtt{in}}}$ with size $H\times W$ and dimension $D_{\mathtt{in}}$, is first flattened to patch tokens $F^{{}^{\prime}}\in$ $\mathbb{R}^{N\times D_{\mathtt{in}}}$, where $N=H\times W$. Then the tokens are mapped to the slots using the dot product attention module. That is, the attention weight can be calculated as $A=\frac{1}{\sqrt{D}}k(F^{{}^{\prime}})\cdot q(S)\in\mathbb{R}^{N\times K}$, and $q$ and $k$ are linear transformations that map the input $F$ and slots $S$ to a common dimension $D$.
@@ 3.3 Action-slot
In this section, we describe the details of our design for learning action-centric representations. We follow [5] to use learnable parameters for slot initializations. We extract image features $F\in\mathbb{R}^{T\times H\times W\times D_{in}}$ from the convolution-based encoder [33, 12, 28, 27] for each frame in a video clip and then flatten them to $F^{{}^{\prime}}\in\mathbb{R}^{N\times D_{\mathtt{in}}}$, where $N=T\times H\times W$. We add learnable 3D spatial-temporal positional embeddings $E\in\mathbb{R}^{T\times H\times W}$ to the tokens $F^{{}^{\prime}}$ [7, 4, 26, 5, 70].
@@
Instead of providing attention guidance for positive classes, which is non-trivial as explained previously, we introduce a regularization term to discourage action slots associated with negative classes (i.e., those activities not present in a video), from attending to any regions. We construct an attention mask $M_{\mathtt{neg}}$ wherein all element values are set to 0. We design a loss function: $L_{\mathtt{neg}}=\sum_{\{c|y_{c}=0\}}\textnormal{BCE}(\bar{A}_{c},M_{\mathtt{neg}})$, where $\bar{A}_{c}$ is the normalized attention output of slot $S_{c}$ for negative class $c$. By regularizing the action slots with negative classes, the competition mechanism makes slots with positive classes more competitive in the important regions. Note that the regularization does not require additional annotations. Finally, the full objective function of Action-slot is:
$L_{\mathtt{all}}=L_{\mathtt{act}}+w_{\mathtt{bg}}L_{\mathtt{bg}}+w_{\mathtt{neg}}L_{\mathtt{neg}}$
, where $w_{\mathtt{bg}}$ and $w_{\mathtt{neg}}$ are the weights to balance two terms (set as 0.5 and 1, respectively in this paper).
@@ 4.1 Datasets
TACO. We construct the TACO dataset in the CARLA simulator [22] to attain a balanced distribution encompassing 64 classes of traffic atomic activities. We leverage three approaches to collect data: auto-pilot, scenario-runner [2], and automatic scenario generation [44]. The TACO dataset comprises 5178 clips. Among them, 1148 are used for testing. We collect image and instance segmentation data with a size of 512 $\times$ 1536 pixels, and subsequently downsample each frame to 256 $\times$ 768 as input to the models.
@@ 4.2 Baselines
Object-aware models. The models [6, 74, 3] first extract object features via the RoI-Align [34] and learn the relations between objects. ORN [6], ARG [74], and OATS [3] uses MLP, Graph Convolutional Network (GCN) [42], and spatical-temporal GCN [52] to enhance object features. OATS also uses an additional graph branch to encode low-level object trajectories. We use Deep OC-SORT [50] to extract trajectories of objects. We follow OATS by setting the number of tracklets as 20 for the input of the models. To encourage the object-aware models [6, 74] to learn the context information, we concatenate object features with the global features, which are obtained by convolving X3D [27] features with a 3D kernel size of 1. These models then output multi-class predictions for each proposal by a linear layer with a channel number of $N_{cl}$ plus one negative class. Note that we do not implement the model proposed in OATS [3] due to the absence of source code.
Slot-based models. The models take features from the last ConvBlock of the encoders [33, 27] as input. Note that we remove the last projection layer of X3D, which makes the size of some slot-based models slightly smaller than X3D. We re-implement existing slot-attention-based methods for videos [41, 24, 5, 81] to use the proposed “allocated slots" for multi-label atomic activity recognition. Specifically, SAVi [41] randomly initialize slots for each forward pass from a learned distribution, while MO [5] and Slot-VPS [81] treat slots as learnable queries [11]. For the slot update mechanism, SAVi and MO recurrently update slots in time, while Slot-VPS uses an extra self-attention [72] to update slots across frames. We then use their slots in the last frame as the input to the classifier.
@@
Method Backbone Seq Pretrain S1 S2 S3 mAP
CSN [71, 3] ResNet152 32 IG65M 12.1 12.6 12.9 12.5
TPN [78, 3] ResNet50 32 ImageNet 11.6 13.3 12.9 12.6
SlowOnly [28, 3] ResNet50 32 ImageNet 11.2 14.7 12.9 13.0
SlowFast [28, 3] ResNet50 32 None 10.8 15.1 14.5 13.5
I3D (NL) [73, 3] ResNet50 32 ImageNet 11.9 15.5 14.0 13.8
I3D [12, 3] ResNet50 32 ImageNet 11.8 14.3 16.8 14.3
ORN [6, 3] ResNet50 32 ImageNet 16.8 13.4 18.1 16.1
ARG [74, 3] Inceptionv3 32 ImageNet 20.2 21.3 19.3 20.3
OATS [3] Inceptionv3 32 ImageNet 24.3 28.6 27.2 26.7
@@
Action-slot (Ours) ResNet50 32 ImageNet 26.6 28.6 30.8 28.6
I3D [12] ResNet50 8 Kinetics-400 21.7 24.6 24.4 23.6
X3D [27] N/A 16 Kinetics-400 30.4 33.2 30.6 31.4
CSN [71] ResNet101 32 Kinetics-400 43.1 47.1 44.3 44.8
SlowFast [28] ResNet50 16 Kinetics-400 36.1 36.6 34.2 35.6
ORN [6] X3D 16 Kinetics-400 19.3 22.5 23.6 21.8
ARG [74] X3D 16 Kinetics-400 24.8 25.9 29.3 26.7
@@
Action-slot (Ours) X3D 16 Kinetics-400 48.1 47.7 48.8 48.2
@@ 4.3 Implementation Details
Table 5: Results of various backbones for Action-slot on TACO.
@@ 4.4 Main Results
OATS. In Table 1, we show experimental comparisons with video-level-based, object-aware, and slot-attention-based methods. In the top group, we use the ImageNet pre-trained ResNet50, following the evaluation protocol designed in OATS [3]. We further compare all methods with advanced visual backbones pretrained from Kinetics-400 [40] and report the corresponding results in the bottom group. In both ImageNet and Kinetics pretrained settings, Action-slot achieves state-of-the-art performance on the OATS dataset. Specifically, Action-slot improves the performance of atomic activity recognition by identifying the region of interest via the proposed slot attention mechanism, without explicit perception guidance.
@@ 4.5 Ablation Study
Figure 6: Visualization of attention maps learned from TACO. Colored masks are from the allocated slots of activity Z4-Z1:K and Z1-Z2:K+. Note that Z4-Z1:K appears twice in the 1st and 5th frames independently.
@@ 6 Conclusions
Acknowledgement. The work is sponsored in part by the Higher Education Sprout Project of the National Yang Ming Chiao Tung University and Ministry of Education (MOE), the Yushan Fellow Program Administrative Support Grant, and the National Science and Technology Council (NSTC) under grants 110-$2$-E-A49-001-MY3, 111-2634-F-002-022-, 113-2923-E-A49 -003 -MY2, Mobile Drive Technology Co., Ltd (MobileDrive), and Industrial Technology Research Institute Mechanical and Mechatronics Systems Lab.
@@ Appendix A Construction of The TACO Dataset
We collect scenarios with the following three approaches:
@@
    Automatic Scenario Generation [44]. We use existing pre-recorded basic scenarios in RiskBench [44] and automatically generate diverse scenarios from a basic scenario. Specifically, we select the pre-recorded basic scenarios from interactive scenarios where ego-vehicle interacts with other risky road users and non-interactive scenarios where ego-vehicle does not interact with any road user. Then we follow the augmentation process proposed in RiskBench [44] to automatically generate diverse scenarios via injecting random road users and changing the weather. The scenarios collected by RiskBench provide more human-like maneuvers and more risky interactions compared to the scenarios collected by the auto-pilot.
  3. (c)
    Scenario Runner [2]. Scenario runner, a scenario collection tool developed by CARLA [22], can explicitly generate a scripted scenario by defining a set of routes for road users. Although auto-pilot and automatic scenario generation can help collect diverse traffic activities, it is difficult to generate some specific classes of atomic activity frequently. For example, Z3-Z2:K+, a group of bicyclists turns left from the opposite roadway of ego-vehicle. Therefore, we leverage scenario runner to explicitly collect the scenarios that are difficult for the auto-pilot method. A scenario starts to collect when ego vehicle reaches a trigger point (i.e., location) predefined in the script. A scenario ends when all scripted actions are accomplished. To further enhance the diversity of scenarios, we randomly spawn road users surrounding the ego vehicle. Note that road users are not guaranteed to be involved in an atomic activity.
For all collection methods, we randomly set the weather and light conditions. Note that we exclude night scenes because of the poor visibility.
@@ Appendix C More Ablation Study of Action-slot
Object-guided vs. Action-slot.
@@
Cross-Attention vs. Slot Attention.